OpenStudy (anonymous):

Solve, using the substitution method. 5x + 2y = 7 y = x – 7 (Points : 1) (7, 0) (3, –4) (0, –7) (–1, –8)

OpenStudy (anonymous):

sub x - 7 in for y 5x + 2(x - 7) = 7 5x + 2x - 14 = 7 5x + 2x = 7 + 14 7x = 21 x = 3 now sub 3 in for x y = x - 7 y = 3 - 7 y = -4 solution is : (3,-4) any questions ?

OpenStudy (anonymous):

how r these type of things so simple 2 u do u think u could help me with it

OpenStudy (anonymous):

yes...I can...what is it that you do not understand ?

OpenStudy (anonymous):

all of it lol Equations inequalities

OpenStudy (anonymous):

inequalities are solved just like equations except, with inequalities, when you divide/multiply by a negative, the inequality sign changes. The main thing is getting the variable terms on one side and everything else on the other side. EX. 2x + 3 = 5x - 4 -- subtract 5x from both sides 2x + 3 - 5x = 5x - 5x - 4 -- combine like terms -3x + 3 = -4 -- now subtract 3 from both sides -3x + 3 - 3 = -4 - 3 -- combine like terms -3x = -7 -- now divide by -3 on both sides -3x/-3 = -7/-3 x = 7/3 you have to make sure that what you do on one side, you have to do on the other side. One thing that might help is, if you want something moved to the other side of the equal sign, you do the opposite of what is showing. Ex. 2x - 3 = 3x + 2 -- if you want to move the -3 to the other side, you ADD 3. 5x + 3 = 2x + 6 -- if you want to move the 3 to the other side, you SUBTRACT 3 same with division and multiplication. 3/7x = 9 -- forget about the x for a sec. You have a problem showing division. The opposite of division is multiplication. So we multiply by 7. (3/7x)/7 = (9)(7) -- the 7's cancel out on the left side, leaving us with : 3x = 63 -- now it shows multiplication...so we do the opposite and divide by 3. 3x/3 = 63/3 -- the 3's on the left side cancel out. x = 63/3 x = 21 Is any of this making sense to you ?

OpenStudy (anonymous):

surprisingly yes lol